OpenStudy (anonymous):

Solve for x. 5x-1/4=6 What value of x makes the equation true?

OpenStudy (anonymous):

@ikram002p

OpenStudy (astrophysics):

\[\frac{ 5x-1 }{ 4 }=6\] like this?

OpenStudy (anonymous):

yes

OpenStudy (astrophysics):

You can cross multiply

OpenStudy (astrophysics):

\[5x-1 = 6(4)\] now solve for x

OpenStudy (anonymous):

-19

OpenStudy (astrophysics):

No, not quite, how did you get that?

OpenStudy (ikram002p):

is the equation like this \(\Large 5x-\frac{1}{4}=6 \) or \(\Large \frac{5x-1}{4}=6 \) ??

OpenStudy (anonymous):

the 2nd one

OpenStudy (ikram002p):

oh then gon on with @Astrophysics

OpenStudy (astrophysics):

When doing algebra, what you do to one side you must also do to the other, and to figure out what operation to do, just do the opposite of what already exists :) So see we have 5x-1 = 20 and we want to get x on its own. So first lets get rid of the -1, to do that as I mentioned above do the opposite operation of it and then do it to both left and right side as follow

OpenStudy (astrophysics):

 now what's next